# CST Anti-UAV: A Thermal Infrared Benchmark for Tiny UAV Tracking in Complex Scenes

### v1

@@ Abstract
††🖂 : Corresponding author.The widespread application of Unmanned Aerial Vehicles (UAVs) has raised serious public safety and privacy concerns, making UAV perception crucial for anti-UAV tasks. However, existing UAV tracking datasets predominantly feature conspicuous objects and lack diversity in scene complexity and attribute representation, limiting their applicability to real-world scenarios. To overcome these limitations, we present the CST Anti-UAV, a new thermal infrared dataset specifically designed for Single Object Tracking (SOT) in Complex Scenes with Tiny UAVs (CST). It contains 220 video sequences with over 240k high-quality bounding box annotations, highlighting two key properties: a significant number of tiny-sized UAV targets and the diverse and complex scenes. To the best of our knowledge, CST Anti-UAV is the first dataset to incorporate complete manual frame-level attribute annotations, enabling precise evaluations under varied challenges. To conduct an in-depth performance analysis for CST Anti-UAV, we evaluate 20 existing SOT methods on the proposed dataset. Experimental results demonstrate that tracking tiny UAVs in complex environments remains a challenge, as the state-of-the-art method achieves only 35.92% state accuracy, much lower than the 67.69% observed on the Anti-UAV410 dataset. These findings underscore the limitations of existing benchmarks and the need for further advancements in UAV tracking research. The CST Anti-UAV benchmark is about to be publicly released, which not only fosters the development of more robust SOT methods but also drives innovation in anti-UAV systems.
@@
| PTB-TIR [35] | 60 | 30k | 9 | 60/60 | - | - | 0 | 1883 | 6325 | 73925 |
| LSOTB-TIR [36] | 1400 | 606k | 12 | 120/1400 | - | - | 0 | 7625 | 30910 | 568294 |
| Anti-UAV [29] | 318 | 293k | 7 | 318/318 | - | - | 858 | 6419 | 88181 | 198317 |
| Anti-UAV410 [26] | 410 | 438k | 6 | 410/410 | 6 | 701k/2628k | 17071 | 126649 | 97942 | 196735 |
@@ 1 Introduction
UAVs have rapidly proliferated due to their low cost and versatile applications in domains such as surveillance, delivery, and imaging [12, 30, 25, 62]. However, incidents of unauthorized or malicious drone flights have infringed on privacy and endangered public safety, raising serious security and privacy concerns [44]. Consequently, there is a growing need for anti-UAV systems capable of real-time detection and tracking of unauthorized UAVs. UAV tracking is a crucial component of such systems, providing the location and trajectory of the target drone for interception or other countermeasures.
Most existing SOT datasets [50, 51, 45, 33] are based on visible light images, leading to unreliable tracking results in low-light or foggy conditions. Although some Thermal Infrared (TIR) datasets [21, 20, 42, 52] have been proposed, they focus on general objects and may not be optimal for tiny UAV tracking. Anti-UAV [29] is the first dataset specifically designed for UAV tracking. Anti-UAV410 [26] was further expanded based on the Anti-UAV dataset by collecting some tiny UAVs and wild scenes. However, the existing datasets for tracking UAVs still have limitations as follows: 1) They lack a sufficient number of tiny UAV targets for effective training, thus limiting real-world applications. 2) The backgrounds of the existing datasets are relatively clean and quite simple, failing to capture the complexity of real-world tracking scenarios. 3) They have incomplete frame-level attribute annotations, hampering thorough evaluations of tracking methods.
To tackle the above problems, we propose a large-scale dataset for tracking UAVs in complex scenes, referred to as CST Anti-UAV. We use popular TIR single-object tracking datasets, including PTB-TIR [35], LSOTB-TIR [36], Anti-UAV [29], and Anti-UAV410 [26], as references to analyze the data statistics of the new CST Anti-UAV dataset. As shown in Table 1, the CST Anti-UAV dataset contains 220 sequences and over 240k high-quality bounding boxes, surpassing existing datasets in three aspects:
@@
To analyze the proposed dataset, we retrained the 20 recent SOT methods on our dataset and conducted extensive evaluations. The experimental results demonstrate that complex tracking scenes and tiny-size objects make the SOT methods less pronounced. For example, SiamDT’s [26] State Accuracy (SA) performance drops from 67.69% on Anti-UAV410 to 35.84% on CST Anti-UAV, while GlobalTrack [27] declines from 66.42% to 35.92%, consistently revealing the challenges brought by the CST Anti-UAV dataset.
@@ 2 Related Work
Single object tracking methods. Based on the latest developments in visual tracking, the tracking methods can be classified into three groups: Transformer-based methods [9, 7, 57, 53, 16, 54, 34, 46, 8] leverage the self-attention mechanism to model long-range dependencies and capture global context. Siamese methods [60, 22, 61, 58, 56, 2, 11, 18, 32] employ twin networks to compare template and search features, balancing speed and accuracy. Other deep learning methods [4, 3, 37, 16, 15, 47, 40, 41] integrate diverse architectures like convolutional neural networks or attention mechanisms to address specific tracking challenges.
Single object tracking benchmarks. TrackingNet [39] is a visible object tracking benchmark dataset based on YouTube-BoundingBoxes [43], containing over 30,000 video clips with 27 target categories. LaSOT [19] is a high-quality benchmark that includes 1,400 visible videos, with an average of around 2,500 frames per video. GOT10K [28] is a visible object tracking dataset that covers 563 types of objects, containing more than 1.5 million manually annotated bounding boxes. PTB-TIR [35] is a thermal infrared pedestrian tracking dataset that contains 60 sequences collected from various devices. LSOTB-TIR [36] is currently the largest TIR object tracking dataset with 1,400 sequences, defining 12 challenge attributes.
Benchmarks for anti-UAV. Although several anti-UAV datasets have been proposed, they are still far from being practically applicable. Anti-UAV [29] is a dataset with visible and infrared dual-mode information, which consists of 318 fully labeled videos, making a significant contribution to maintaining airspace safety. However, most objects in the dataset are relatively large, and the backgrounds are clean and simple. DUT Anti-UAV [63] is divided into two tasks: detection and tracking. It contains 10,000 images for the detection task and 20 sequences for the tracking task. However, the small number of tracking sequences is insufficient for training tracking methods. Anti-UAV410 [26] extends Anti-UAV to track drones in wild scenes. It has collected 150k drone objects and merged the infrared portion of the Anti-UAV dataset, ultimately containing 410 sequences with 438k objects. The collected data includes some tiny objects and wild scenes. Unfortunately, the number of tiny objects is too limited, and the scenes are predominantly in wilderness areas, lacking truly complex urban areas with dynamic distractors involving pedestrians and vehicles. The singular attribute of sequences fails to capture the complexity of real tracking scenes. Furthermore, the frame-level annotations are incomplete, hindering comprehensive evaluation of tracker performance across various attributes.
@@ 3.1 Data Acquisition and Analysis
Large-scale sequences with high diversity. Existing datasets are [29] typically recorded through a rotating platform equipped with a static camera, the acquisition scenes are limited only to a few selected. Therefore, we simultaneously adopt rotating platforms and professional drones equipped with infrared cameras (25 fps, 640 512) for data collection to ensure diversity. We carefully captured 220 sequences, comprising over 240k high-quality bounding boxes, and sequence lengths range from 600 to 2,062 frames (average 1,132 frames), covering both short-term and extended tracking scenarios. Furthermore, we classified UAV objects into four size categories, with their proportional distribution illustrated in the inner ring of Figure 2 (a). Notably, small and tiny objects are the domain, reflecting real-world tracking scenarios where UAVs are typically inconspicuous. Tiny objects are more difficult to track, while larger objects require fewer samples to achieve comparable performance. A key highlight is that our dataset contains 4.5 times more tiny objects than existing datasets. We further analyzed object size distribution compared to the Anti-UAV [29] and Anti-UAV410 [26] datasets, as shown in Figure 2 (c).
@@
Coarse-to-fine attribute annotation. The challenges in our dataset can be summarized as six attributes, namely Occlusion (OC), Out-of-View (OV), Scale Variation (SV), Thermal Crossover (TC), Dynamic Background Clutter (DBC), and Complex Dynamic Background (CDB), the detailed meanings of each attribute are as shown in Table 2. In previous datasets, the backgrounds were mostly clean and simple. However, in real-world operational environments including aerial surveying and traffic monitoring, backgrounds inevitably contain complex dynamic objects like pedestrians and vehicles. We propose the novel concept of CDB which incorporates these dynamic objects. In comparison with existing anti-UAV datasets [29, 26] on sequence-level attributes, as shown in Figure 2 (b). The number of almost all attributes is significantly higher than in existing datasets. Notably, the CST Anti-UAV dataset contains 184 sequences with more than three attributes, compared to only 60 sequences in the Anti-UAV410 dataset. Furthermore, existing infrared tracking datasets only annotate attributes at the sequence level. In fact, only a few frames in a sequence may correspond to a specific challenge, roughly evaluating multiple different challenges across an entire sequence leads to flawed analysis results. We first introduce full frame-level attribute annotations. As shown in Figure 2 (d), we show the number of frames for each challenge attribute.
@@ 3.2 High-quality Human Annotation
After collecting all the videos for the CST Anti-UAV dataset, we reviewed and edited them to reduce the number of inappropriate frames, resulting in over 240k frames. To ensure the quality of the annotations, we invested over 2,000 hours in the annotation process, which was divided into two tracks: 1) Bounding box annotation. Unlike existing datasets [29, 36] that rely on sparse annotation or algorithmic pre-labeling with manual correction, our annotation team manually obtained high-quality bounding boxes for each frame. Our verification team conducted continuous frame-by-frame checks until no issues remained. 2) Attribute annotation. To ensure consistency in evaluating the six proposed challenge attributes across frames, we assigned a single expert annotator to handle all attribute labeling, given that different annotators may exhibit varying tendencies. In total, we annotated 1440k attributes across all frames and provided 220 6 sequence-level attribute annotations.
@@ 3.3 Dataset Splitting
| GlobalTrack [27] | AAAI20 | 66.42 | 85.25 | 65.11 | 35.92 | 58.72 | 35.38 |
| KYS [4] | ECCV20 | 45.12 | 61.31 | 44.30 | 26.28 | 40.20 | 25.93 |
| PrDiMP [17] | CVPR20 | 59.31 | 76.77 | 59.09 | 29.76 | 42.55 | 29.28 |
| Stark [57] | ICCV21 | 57.09 | 76.64 | 55.99 | 23.27 | 40.54 | 23.10 |
| Stark-ST [57] | ICCV21 | 58.32 | 78.10 | 57.22 | 25.63 | 44.99 | 25.44 |
| OSTrack-256 [59] | ECCV22 | 52.47 | 70.39 | 51.44 | 26.34 | 46.95 | 26.16 |
| OSTrack-384 [59] | ECCV22 | 57.41 | 76.73 | 56.29 | 25.25 | 45.99 | 25.10 |
| AiATrack [23] | ECCV22 | 58.31 | 78.15 | 57.19 | 26.83 | 47.65 | 26.64 |
| ToMP [38] | CVPR22 | 56.44 | 73.70 | 55.28 | 22.84 | 39.81 | 22.66 |
| Mixformer [13] | CVPR22 | 58.49 | 78.07 | 57.35 | 26.46 | 47.14 | 26.19 |
| Mixformer-V2 [14] | NIPS23 | 61.68 | 81.41 | 60.50 | 29.39 | 52.37 | 29.09 |
| DropTrack [49] | CVPR23 | 61.00 | 81.11 | 59.81 | 26.89 | 49.08 | 26.65 |
| ARTrack-256 [48] | CVPR23 | 48.14 | 63.75 | 47.17 | 25.14 | 45.39 | 24.96 |
| ARTrack-384 [48] | CVPR23 | 52.63 | 69.51 | 51.54 | 25.19 | 46.56 | 25.04 |
| SeqTrack [10] | CVPR23 | 50.50 | 67.26 | 49.49 | 26.53 | 47.67 | 26.38 |
| GRM [24] | CVPR23 | 50.79 | 68.36 | 49.80 | 26.96 | 48.70 | 26.80 |
| ROMTrack [6] | ICCV23 | 55.08 | 72.62 | 53.98 | 27.43 | 48.20 | 27.23 |
| HIPTrack [5] | CVPR24 | 57.78 | 77.59 | 56.64 | 25.62 | 47.48 | 25.42 |
| AQATrack [55] | CVPR24 | 56.64 | 75.33 | 55.52 | 27.60 | 50.43 | 27.45 |
| ARTrackV2 [1] | CVPR24 | 52.51 | 70.41 | 51.49 | 22.14 | 39.73 | 22.03 |
| SiamDT* [26] | PAMI24 | 67.69 | 87.63 | 66.34 | 35.84 | 57.96 | 35.28 |
| Refocus-TIR* [31] | TNNLS24 | 58.24 | 76.31 | 57.09 | 26.51 | 47.49 | 26.26 |

### v2

@@ Abstract
††🖂 : Corresponding author.The widespread application of Unmanned Aerial Vehicles (UAVs) has raised serious public safety and privacy concerns, making UAV perception crucial for anti-UAV tasks. However, existing UAV tracking datasets predominantly feature conspicuous objects and lack diversity in scene complexity and attribute representation, limiting their applicability to real-world scenarios. To overcome these limitations, we present the CST Anti-UAV, a new thermal infrared dataset specifically designed for Single Object Tracking (SOT) in Complex Scenes with Tiny UAVs (CST). It contains 220 video sequences with over 240k high-quality bounding box annotations, highlighting two key properties: a significant number of tiny-sized UAV targets and the diverse and complex scenes. To the best of our knowledge, CST Anti-UAV is the first dataset to incorporate complete manual frame-level attribute annotations, enabling precise evaluations under varied challenges. To conduct an in-depth performance analysis for CST Anti-UAV, we evaluate 20 existing SOT methods on the proposed dataset. Experimental results demonstrate that tracking tiny UAVs in complex environments remains a challenge, as the state-of-the-art method achieves only 35.92% state accuracy, much lower than the 67.69% observed on the Anti-UAV410 dataset. These findings underscore the limitations of existing benchmarks and the need for further advancements in UAV tracking research. The CST-Anti-UAV benchmark is now publicly available at https://github.com/PCwenyue/CST-Anti-UAV.
@@
| PTB-TIR [ptb-tir] | 60 | 30k | 9 | 60/60 | - | - | 0 | 1883 | 6325 | 73925 |
| LSOTB-TIR [lsotb-tir] | 1400 | 606k | 12 | 120/1400 | - | - | 0 | 7625 | 30910 | 568294 |
| Anti-UAV [antiuav] | 318 | 293k | 7 | 318/318 | - | - | 858 | 6419 | 88181 | 198317 |
| Anti-UAV410 [antiuav410] | 410 | 438k | 6 | 410/410 | 6 | 701k/2628k | 17071 | 126649 | 97942 | 196735 |
@@ 1 Introduction
UAVs have rapidly proliferated due to their low cost and versatile applications in domains such as surveillance, delivery, and imaging [tuav, uav3, uav4, uav5]. However, incidents of unauthorized or malicious drone flights have infringed on privacy and endangered public safety, raising serious security and privacy concerns [uav2]. Consequently, there is a growing need for anti-UAV systems capable of real-time detection and tracking of unauthorized UAVs. UAV tracking is a crucial component of such systems, providing the location and trajectory of the target drone for interception or other countermeasures.
Most existing SOT datasets [otb50, otb100, alov, tc128] are based on visible light images, leading to unreliable tracking results in low-light or foggy conditions. Although some Thermal Infrared (TIR) datasets [vottir2016, vottir2015, pdt-atv, bu-tiv] have been proposed, they focus on general objects and may not be optimal for tiny UAV tracking. Anti-UAV [antiuav] is the first dataset specifically designed for UAV tracking. Anti-UAV410 [antiuav410] was further expanded based on the Anti-UAV dataset by collecting some tiny UAVs and wild scenes. However, the existing datasets for tracking UAVs still have limitations as follows: 1) They lack a sufficient number of tiny UAV targets for effective training, thus limiting real-world applications. 2) The backgrounds of the existing datasets are relatively clean and quite simple, failing to capture the complexity of real-world tracking scenarios. 3) They have incomplete frame-level attribute annotations, hampering thorough evaluations of tracking methods.
To tackle the above problems, we propose a large-scale dataset for tracking UAVs in complex scenes, referred to as CST Anti-UAV. We use popular TIR single-object tracking datasets, including PTB-TIR [ptb-tir], LSOTB-TIR [lsotb-tir], Anti-UAV [antiuav], and Anti-UAV410 [antiuav410], as references to analyze the data statistics of the new CST Anti-UAV dataset. As shown in Table 1, the CST Anti-UAV dataset contains 220 sequences and over 240k high-quality bounding boxes, surpassing existing datasets in three aspects:
@@
To analyze the proposed dataset, we retrained the 20 recent SOT methods on our dataset and conducted extensive evaluations. The experimental results demonstrate that complex tracking scenes and tiny-size objects make the SOT methods less pronounced. For example, SiamDT’s [antiuav410] State Accuracy (SA) performance drops from 67.69% on Anti-UAV410 to 35.84% on CST Anti-UAV, while GlobalTrack [globaltrack] declines from 66.42% to 35.92%, consistently revealing the challenges brought by the CST Anti-UAV dataset.
@@ 2 Related Work
Single object tracking methods. Based on the latest developments in visual tracking, the tracking methods can be classified into three groups: Transformer-based methods [transt, hift, stark, sbt, atom, supersbt, swintrack, cswintt, simtrack] leverage the self-attention mechanism to model long-range dependencies and capture global context. Siamese methods [siamattn, stmtrack, ocean, alpha, siamfc++, siamfc, siamban, cgacd, siamrpnpp] employ twin networks to compare template and search features, balancing speed and accuracy. Other deep learning methods [kys, dimp, keeptrack, atom, eco, udt, mdnet, metasdnet] integrate diverse architectures like convolutional neural networks or attention mechanisms to address specific tracking challenges.
Single object tracking benchmarks. TrackingNet [tracking] is a visible object tracking benchmark dataset based on YouTube-BoundingBoxes [ytbb], containing over 30,000 video clips with 27 target categories. LaSOT [lasot] is a high-quality benchmark that includes 1,400 visible videos, with an average of around 2,500 frames per video. GOT10K [got10k] is a visible object tracking dataset that covers 563 types of objects, containing more than 1.5 million manually annotated bounding boxes. PTB-TIR [ptb-tir] is a thermal infrared pedestrian tracking dataset that contains 60 sequences collected from various devices. LSOTB-TIR [lsotb-tir] is currently the largest TIR object tracking dataset with 1,400 sequences, defining 12 challenge attributes.
Benchmarks for anti-UAV. Although several anti-UAV datasets have been proposed, they are still far from being practically applicable. Anti-UAV [antiuav] is a dataset with visible and infrared dual-mode information, which consists of 318 fully labeled videos, making a significant contribution to maintaining airspace safety. However, most objects in the dataset are relatively large, and the backgrounds are clean and simple. DUT Anti-UAV [dutantiuav] is divided into two tasks: detection and tracking. It contains 10,000 images for the detection task and 20 sequences for the tracking task. However, the small number of tracking sequences is insufficient for training tracking methods. Anti-UAV410 [antiuav410] extends Anti-UAV to track drones in wild scenes. It has collected 150k drone objects and merged the infrared portion of the Anti-UAV dataset, ultimately containing 410 sequences with 438k objects. The collected data includes some tiny objects and wild scenes. Unfortunately, the number of tiny objects is too limited, and the scenes are predominantly in wilderness areas, lacking truly complex urban areas with dynamic distractors involving pedestrians and vehicles. The singular attribute of sequences fails to capture the complexity of real tracking scenes. Furthermore, the frame-level annotations are incomplete, hindering comprehensive evaluation of tracker performance across various attributes.
@@ 3.1 Data Acquisition and Analysis
Large-scale sequences with high diversity. Existing datasets are [antiuav] typically recorded through a rotating platform equipped with a static camera, the acquisition scenes are limited only to a few selected. Therefore, we simultaneously adopt rotating platforms and professional drones equipped with infrared cameras (25 fps, 640 512) for data collection to ensure diversity. We carefully captured 220 sequences, comprising over 240k high-quality bounding boxes, and sequence lengths range from 600 to 2,062 frames (average 1,132 frames), covering both short-term and extended tracking scenarios. Furthermore, we classified UAV objects into four size categories, with their proportional distribution illustrated in the inner ring of Figure 2 (a). Notably, small and tiny objects are the domain, reflecting real-world tracking scenarios where UAVs are typically inconspicuous. Tiny objects are more difficult to track, while larger objects require fewer samples to achieve comparable performance. A key highlight is that our dataset contains 4.5 times more tiny objects than existing datasets. We further analyzed object size distribution compared to the Anti-UAV [antiuav] and Anti-UAV410 [antiuav410] datasets, as shown in Figure 2 (c).
@@
Coarse-to-fine attribute annotation. The challenges in our dataset can be summarized as six attributes, namely Occlusion (OC), Out-of-View (OV), Scale Variation (SV), Thermal Crossover (TC), Dynamic Background Clutter (DBC), and Complex Dynamic Background (CDB), the detailed meanings of each attribute are as shown in Table 2. In previous datasets, the backgrounds were mostly clean and simple. However, in real-world operational environments including aerial surveying and traffic monitoring, backgrounds inevitably contain complex dynamic objects like pedestrians and vehicles. We propose the novel concept of CDB which incorporates these dynamic objects. In comparison with existing anti-UAV datasets [antiuav, antiuav410] on sequence-level attributes, as shown in Figure 2 (b). The number of almost all attributes is significantly higher than in existing datasets. Notably, the CST Anti-UAV dataset contains 184 sequences with more than three attributes, compared to only 60 sequences in the Anti-UAV410 dataset. Furthermore, existing infrared tracking datasets only annotate attributes at the sequence level. In fact, only a few frames in a sequence may correspond to a specific challenge, roughly evaluating multiple different challenges across an entire sequence leads to flawed analysis results. We first introduce full frame-level attribute annotations. As shown in Figure 2 (d), we show the number of frames for each challenge attribute.
@@ 3.2 High-quality Human Annotation
After collecting all the videos for the CST Anti-UAV dataset, we reviewed and edited them to reduce the number of inappropriate frames, resulting in over 240k frames. To ensure the quality of the annotations, we invested over 2,000 hours in the annotation process, which was divided into two tracks: 1) Bounding box annotation. Unlike existing datasets [antiuav, lsotb-tir] that rely on sparse annotation or algorithmic pre-labeling with manual correction, our annotation team manually obtained high-quality bounding boxes for each frame. Our verification team conducted continuous frame-by-frame checks until no issues remained. 2) Attribute annotation. To ensure consistency in evaluating the six proposed challenge attributes across frames, we assigned a single expert annotator to handle all attribute labeling, given that different annotators may exhibit varying tendencies. In total, we annotated 1440k attributes across all frames and provided 220 6 sequence-level attribute annotations.
@@ 3.3 Dataset Splitting
| GlobalTrack [globaltrack] | AAAI20 | 66.42 | 85.25 | 65.11 | 35.92 | 58.72 | 35.38 |
| KYS [kys] | ECCV20 | 45.12 | 61.31 | 44.30 | 26.28 | 40.20 | 25.93 |
| PrDiMP [prdimp] | CVPR20 | 59.31 | 76.77 | 59.09 | 29.76 | 42.55 | 29.28 |
| Stark [stark] | ICCV21 | 57.09 | 76.64 | 55.99 | 23.27 | 40.54 | 23.10 |
| Stark-ST [stark] | ICCV21 | 58.32 | 78.10 | 57.22 | 25.63 | 44.99 | 25.44 |
| OSTrack-256 [ostrack] | ECCV22 | 52.47 | 70.39 | 51.44 | 26.34 | 46.95 | 26.16 |
| OSTrack-384 [ostrack] | ECCV22 | 57.41 | 76.73 | 56.29 | 25.25 | 45.99 | 25.10 |
| AiATrack [aiatrack] | ECCV22 | 58.31 | 78.15 | 57.19 | 26.83 | 47.65 | 26.64 |
| ToMP [tomp] | CVPR22 | 56.44 | 73.70 | 55.28 | 22.84 | 39.81 | 22.66 |
| Mixformer [mixformer] | CVPR22 | 58.49 | 78.07 | 57.35 | 26.46 | 47.14 | 26.19 |
| Mixformer-V2 [mixformerv2] | NIPS23 | 61.68 | 81.41 | 60.50 | 29.39 | 52.37 | 29.09 |
| DropTrack [droptrack] | CVPR23 | 61.00 | 81.11 | 59.81 | 26.89 | 49.08 | 26.65 |
| ARTrack-256 [artrack] | CVPR23 | 48.14 | 63.75 | 47.17 | 25.14 | 45.39 | 24.96 |
| ARTrack-384 [artrack] | CVPR23 | 52.63 | 69.51 | 51.54 | 25.19 | 46.56 | 25.04 |
| SeqTrack [seqtrack] | CVPR23 | 50.50 | 67.26 | 49.49 | 26.53 | 47.67 | 26.38 |
| GRM [grm] | CVPR23 | 50.79 | 68.36 | 49.80 | 26.96 | 48.70 | 26.80 |
| ROMTrack [romtrack] | ICCV23 | 55.08 | 72.62 | 53.98 | 27.43 | 48.20 | 27.23 |
| HIPTrack [hiptrack] | CVPR24 | 57.78 | 77.59 | 56.64 | 25.62 | 47.48 | 25.42 |
| AQATrack [aqatrack] | CVPR24 | 56.64 | 75.33 | 55.52 | 27.60 | 50.43 | 27.45 |
| ARTrackV2 [artrackv2] | CVPR24 | 52.51 | 70.41 | 51.49 | 22.14 | 39.73 | 22.03 |
| SiamDT* [antiuav410] | PAMI24 | 67.69 | 87.63 | 66.34 | 35.84 | 57.96 | 35.28 |
| Refocus-TIR* [refocus] | TNNLS24 | 58.24 | 76.31 | 57.09 | 26.51 | 47.49 | 26.26 |